What's the present value of the right to receive five equal payments (ordinary annuity) of $750 per period, discounted at a rate of 12 percent per period?
Answer #1

PV = PMT x ((1 - (1 / (1 + r) ^ n)) / r)

  = 750*((1-(1/1.12)^5)) / 0.12

PV = $2,703.58

OR

Year Annuity(a) PV Factor @12%(b) Present Value(a*b)
1 750 1/(1.12^1) = 0.8929 669.643
2 750 1/(1.12^2) = 0.7972 597.895
3 750 1/(1.12^3) = 0.7118 533.835
4 750 1/(1.12^4) = 0.6355 476.639
5 750 1/(1.12^5) = 0.5674 425.570

Total 2,703.582